what's the inverse function of 5 to the power of x
Answer by lwsshak3(11628)

what's the inverse function of 5 to the power of x
**
y=5^x (exponential form)
To find inverse:
swap x & y
x=5^y
take log base5 of both sides
log(base5)(x)=log(base5)(5^y)
log(base5)(x)=y (log of base raised to a power=power)
or
y=log(base5)(x) (inverse of y=5^x)
note: for any base, the log function is the inverse of the exponential form